OpenStudy (anonymous):

Greg has a piece of rope that is 18 ft long. He wants to cut it into two pieces so that the longer piece is 2 ft less than 3 times the length of the shorter piece. How long will each piece of rope be?

OpenStudy (danjs):

Longer piece = 3(Shorter piece) - 2

OpenStudy (danjs):

total = shorter + longer

OpenStudy (danjs):

2 equations and 2 unknowns

OpenStudy (anonymous):

thxs a lot Danjs

OpenStudy (danjs):

18 = L + S and L = 3*S - 2

OpenStudy (danjs):

can you do that from there?

OpenStudy (danjs):

18 = 3*s -2 + S 18 + 2 = 4*s 20/4 = S

OpenStudy (watchmath):

Let the shortest part is \(x\). Then \(x+(3x-2)=18\)

OpenStudy (danjs):

shorter is 5 ft long

OpenStudy (danjs):

Longer is 3(shorter) - 2 L = 3(5) - 2 = 15 - 2 = 13 ft Longer is 13 ft

OpenStudy (danjs):

you good on that?

OpenStudy (anonymous):

yes thx

OpenStudy (danjs):

no prob... yw